OpenStudy (anonymous):

37.( I'm a parent and been out of school way too long. I'm having a hard time helping my daughter with is. I don't like to ask for help but would some please explain it to me so I can help her. A painter spends 590 minutes painting a house one day. The painter makes $33.25 per hour. a. Use mixed numbers to write the time in hours. Show your work. b. Estimate the amount the painter is paid. Explain how you found your estimate.

hero (hero):

33 and 1/4 dollars per hour 590/60 * 33.25 = pay

hero (hero):

(590/60)*33.25 = pay 9.63 * 33.25 = 326.95 Almost 327 dollars in pay

OpenStudy (asnaseer):

to convert 590 minutes to hours, 1st divide it by 60 and throw away any remainders this gives us 9 hours next convert the 9 hours back into minutes by multiplying it by 60, so 9 hours = 9 * 60 = 540 minutes then take this away from the orignal time in minutes (590) to get 590 - 540 = 50 minutes so painters time is 9 hours and 50 minutes to estimate how much he got paid, we need to multiply the numbers of hours he worked by his hourly rate. he worked 9 hours and 50 minutes, which is "roughly" 10 hours. so we can estimate his cost as 10 * 33.25 = $332.5

OpenStudy (anonymous):

Thank you all for helping an ol' mom out that is not the greatest with this math...

OpenStudy (asnaseer):

glad to help :-)